# The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profiles

### v1

@@ Abstract
pubyear: 2025pagerange: The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profilesReferences
@@ 1 Introduction
The precise definition of a galaxy cluster’s outer boundary is fundamental to studies of cluster formation, evolution, and cosmology (e.g. Kravtsov & Borgani, 2012; Morandi & Sun, 2016; Haggar et al., 2024). Traditional definitions — deduced from friends-of-friends (FoF) linking lengths or spherical overdensity radii —play a key role in determining cluster masses and membership (e.g. Eke et al., 2004; Lukić et al., 2009), yet these are choices driven primarily by what is practicable with data rather than what is most physically meaningful (e.g. White, 2001; Old et al., 2014, 2015). Establishing a boundary that is physically meaningful and practicably measurable via observations remains a major challenge.
@@
A promising physically motivated boundary is provided by the splashback radius, which marks the apocentre of recently accreted material orbiting within the cluster potential (e.g. Diemer & Kravtsov 2014; Adhikari et al. 2014; More et al. 2015; Diemer 2022). This radius delineates the transition between orbiting and infalling matter, and manifests as a steepening in the halo density profile (which, we note, has historically—and erroneously—been conflated with the splashback radius itself). The radius of this steepest slope is of particular interest because there is now considerable evidence that this steepening has been measured in observations of real galaxy clusters (e.g. More et al. 2016; Baxter et al. 2017; Nishizawa et al. 2018; Chang et al. 2018a; Contigiani et al. 2019; Xu et al. 2024). This is exciting because there is good reason to expect that splashback radius offers a direct probe of halo assembly history and mass accretion rates, thereby connecting cluster outskirts to cosmological growth (More et al., 2015; Deason et al., 2021; Dacunha et al., 2025).
@@
Observational detections of splashback features have been reported using satellite galaxy counts (e.g. More et al. 2016; Baxter et al. 2017; Nishizawa et al. 2018), weak gravitational lensing (e.g. Chang et al. 2018a; Contigiani et al. 2019; Xu et al. 2024), and the Sunyaev-Zel’dovich (SZ) effect (e.g. Shin et al., 2019; Zürcher & More, 2019; Adhikari et al., 2021) However, these come with their own unique limitations. For example, in the case of satellite galaxy counts, accurate measurement of splashback features depends on the properties of satellite tracers in a given cluster (e.g. O’Neil et al., 2022), and accounting for the complexity of their kinematics (e.g. Aung et al., 2021) and structure of the local cosmic web (e.g. Lebeau et al., 2024; Sun et al., 2025). Satellite galaxies are also biased tracers of the cluster potential, with the most massive tending to accumulate more centrally due to dynamical friction (Ludlow et al., 2009; Adhikari et al., 2016). In the case of weak lensing, individual systems may have only a small number of lensed objects.
@@
A promising alternative probe is the diffuse intra-cluster light (ICL) that is formed from the remnants of disrupted galaxies. This material is a potentially powerful tracer of cluster outskirts (Montes & Trujillo, 2019; Contini, 2021), and work by Deason et al. (2021) has shown a direct correspondence between the caustics in the total stellar and dark matter density, indicating a dynamical link between these components. However, there remain important questions about the detailed nature of the stellar splashback structure, its relation to the underlying dark matter, and its observational accessibility, which we seek to answer in this paper.
@@
A deeper understanding of the stellar splashback feature is crucial for a number reasons. The distribution of stars in a cluster reflect physical processes that affect satellite galaxies and their dark matter halos (such as tidal stripping and dynamical friction) and so their distribution encodes information about the interplay between baryons and dark matter in the cluster outskirts (e.g. Brown et al., 2024), which, when combined with splashback and shock structures in the stars and gas (e.g. Deason et al., 2021; Towler et al., 2024; Zhang et al., 2025) can be used to infer dynamical history. Similarly, projected stellar density profiles provide a complementary means of constraining cluster growth histories (e.g. Caminha et al., 2017; Montes & Trujillo, 2018; Kluge et al., 2025; Joshi et al., 2025). This is especially valuable when supplemented by weak lensing and SZ effect measurements; combining different datasets, affected by projection in different way, allow for a more robust inference of splashback features (e.g. Adhikari et al., 2021).
@@
In this paper, we draw upon the statistical sample of massive galaxy clusters from The Three Hundred collaboration’s hydrodynamical galaxy formation simulation suite (cf. Cui et al., 2018, 2022) to study the relationship between splashback features in the dark matter and stellar components. This is a mass complete sample of clusters drawn from a 1 $h^{-1}$ Gpc box, which have a diversity of assembly histories and larger-scale environments. These simulations self-consistently model the physics of galaxy formation, including star formation, black hole growth, supernovae, and AGN feedback. By investigating the stellar splashback structure and its relation to dark matter in hydrodynamical simulations, our aim is to establish a physically grounded framework for interpreting stellar density profiles as tracers of cluster assembly. This approach promises to enhance the utility of stellar observables—such as the ICL—in constraining cluster boundaries and mass accretion histories, thereby broadening the toolkit available for cluster cosmology.
@@ 2 The data
For this work we use data from the Three Hundred Project, a suite of hydrodynamical re-simulations of the most massive 324 galaxy cluster halos in the dark matter-only MultiDark-Planck 2 (MDPL2) simulation (cf. Klypin et al., 2016), a $1h^{-1}$ Gpc box on a side. These clusters have virial masses in the range $6.4\times 10^{14}h^{-1}M_{\odot}\lesssim M_{200}\lesssim 2.6\times 10^{15}h^{-1}M_{\odot}$, where $M_{200}$ is the mass corresponding to an overdensity criterion of 200 times the critical density at that epoch. Each resimulated region extends 15 $h^{-1}$Mpc from the cluster centre at zz=0, corresponding to several virial radii.
In this work we use hydrodynamical simulations run with the galaxy formation model simba (cf. Davé et al., 2019), implemented in the gizmo code (cf. Hopkins, 2015). The gravity solver used in the model derives from gadget (cf. Springel, 2005). gizmo uses a moving mesh scheme to solve hydrodynamics. This overcomes the limitations of classic SPH and captures fluid mixing and instabilities. Galaxy formation processes (e.g. radiative cooling, star formation and feedback, black hole formation and growth, multiple modes of black hole feedback) are modeled using a variant of simba (cf. Davé et al., 2019) that has been calibrated for cluster scales (cf. Cui et al., 2022). gizmo was run on regions from both the $3840^{3}$- and $7680^{3}$-particle MDPL2 boxes, which we refer to as GIZMO_3k and GIZMO_7k, respectively. In this work we mainly present results from the high resolution run, but we verify that the lower resolution run gives identical results.
@@
Data are saved in 129 snapshots from $z=16.98$ to $z=0$. For each cluster, we use group catalogues constructed with the AHF halo finder (cf. Knollmann & Knebe, 2009), which includes information about the stellar and gas content of the main halo and its substructures.
@@ 3.1 Orbiting-infalling decomposition
To identify the splashback of stars and dark matter in the simulated clusters, we classify particles as “orbiting” or “infalling” depending on if the particle has passed through at least one pericenter about the center of mass of the cluster. This is done by recording a change in the sign of a particle’s physical radial velocity (peculiar plus Hubble flow) from negative to positive. To filter out radial velocity sign flips due to orbits within subhalos, we also require that particles have traversed an angle of at least $\pi/2$ around the halo center since entering the cluster region. Any particle that has not had such a sign change is regarded as “infalling”. Though this classification technically considers positive radial velocity particles beyond the turn-around radius as “infalling”, we limit our analysis of the density profiles to radii safely within this boundary.
The result of the orbiting-infalling decomposition for the most massive cluster in the GIZMO_3k run is shown in Figure 1. The top row shows the $x$-$y$ projection of the dark matter (left) and stars (right), while the bottom row is their radial phase space (with Hubble flow added). The phase space shows a clear separation between the infalling stream and the orbiting material, with the infalling material abruptly transitioning into orbiting material at the point of pericenter, $v_{\rm r}=0$. The orbiting particles at the largest radii still have some positive radial velocity, meaning they are still moving outwards and have not quite “splashed back”. Visually, the distribution of the dark matter is well traced by the stellar distribution, and both the orbiting dark matter and orbiting stars extend out slightly beyond the virial radius $R_{\rm vir}$ (Bryan & Norman 1998 definition). See e.g. Diemer (2017); Bakels et al. (2021) for discussion on the extent of the orbits of halo material.
@@
Figure 1: Cluster 001 from the GIZMO_3k run, decomposed into orbiting (coloured) and infalling (white) particles. The top row shows the $x$-$y$ projection of the dark matter (left) and stars (right), while the bottom row shows their radial phase space. The virial radius (see Bryan & Norman 1998 for definition) is marked.
@@
Figure 2: The density profiles of the orbiting and infalling components, fit using models 1 - 3. The BCG component is confined to the inner regions, while the stellar “halo” takes on a similar shape to the orbiting profile of the dark matter. The logarithmic slope profiles are shown in the bottom panels.
@@ 3.3 Profile fitting functions
Figure 3: Comparison between $r_{\rm t}$ and $\beta$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The cyan squares are from fitting the $\Gamma$-binned median profiles.
@@
We fit the profiles using least-squares minimization in the logarithm of the density. The radial bins are weighted equally and the fitting is done between $4\,R_{200{\rm m}}$ and the convergence radius predicted by Power et al. (2003) and Ludlow et al. (2019) for the dark matter-only case. The results of fitting the models 1 - 3 to the orbiting and infalling profiles of an individual cluster (cluster 001) is shown in Figure 2.
@@
where $a$ is the scale factor and $\Delta$ represents a change over one dynamical time, defined as the crossing time $2\,R_{200{\rm m}}/V_{200{\rm m}}$. At $z=0$, one dynamical time in the past corresponds to $a\approx 0.769$. Here, as with everywhere else in this work, $M_{200{\rm m}}$ is calculated using the total density (stars, gas, and dark matter). Progenitor halos are identified as those that contain the largest fraction of the most-bound particles of the $z=0$ halos. We choose six equal-width bins between $\Gamma=0$ and $\Gamma=6$ and calculate the median logarithmic density profile in each bin, in units of $r/R_{200{\rm m}}$. The first four bins contain between 51 and 102 clusters, while the remaining two contain 17 and 10.
@@ 4.1 Fit parameter relations
Figure 4: Comparison between the radius of steepest slope, $r_{\rm steep}$ and $r_{\rm t}$ for the dark matter and stars. The points are colored by accretion rate (Equation 4). The cyan squares are from $\Gamma$-binned median profiles. The black dashed lines are linear fits given by $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.41$ for the dark matter and $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.37$ for the stars.
Figure 5: $r_{\rm steep}$ and $r_{\rm t}$ as a function of accretion rate (Equation 4). The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), while in the bottom row it is the same relation rescaled using the best-fit $r_{\rm t}$-$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025).
@@
The truncation radius is a more direct measure of the size of the splashback boundary (technically the orbiting boundary, since the outermost particles may not have reached apocenter yet) than the radius of steepest slope, which is determined in part by the infalling material. In perfectly spherical collapse, the truncation of the orbiting profile is infinitely sharp and the two radii coincide. However, in realistic simulations, mass is accreted anisotropically and particle orbits can take on a variety of shapes, which has the effect of smoothing out the edge of the orbiting profile. In this case, the radius of steepest slope is not simply related to the edge of the orbiting material, but is rather the point at which the infalling profile becomes dominant enough to strongly influence the total slope. In Figure 4 we plot $r_{\rm steep}$ measured from the total profiles against $r_{\rm t}$. Again, the points are coloured by $\Gamma$ and show the gradient expected from the now well-known $r_{\rm steep}$-$\Gamma$ relation. The cyan points are measurements of $r_{\rm steep}$ from the $\Gamma$-binned median total profile. The $r_{\rm t}$-$r_{\rm steep}$ relation seems to be described well by a line with slope of $\sim 0.83$ and $r_{\rm steep}(r_{\rm t}=0)/R_{200{\rm m}}\approx 0.41$ for the dark matter and $0.37$ for the stars (black dashed lines). On average, the radius of steepest slope is $\sim 0.25\,R_{200{\rm m}}$ larger than the truncation radius.
@@
The gradient with accretion rate seen in the radii in Figures 3 and 4 reflect the now well-established relationship between the splashback radius and accretion rate. This relationship was first reported by Diemer & Kravtsov (2014) for the radius of steepest slope in dark matter-only cosmological simulations, while its theoretical basis was established by Adhikari et al. (2014) using a model based on spherical collapse. The relation was shown by More et al. (2015) to be well-described by an exponential of the form $r_{\rm steep}/R_{200{\rm m}}=a+be^{-\Gamma/c}$, and this same function was shown by Deason et al. (2021) to hold for the stellar $r_{\rm steep}$ of 34 C-EAGLE clusters. Recently, Diemer (2025) showed that the relationship between $r_{\rm t}$ and $\Gamma$ for stacked halos is also well-described by an exponential that has remarkably little dependence on halo mass or cosmology.
@@
Figure 7: Comparison between 3D and projected $r_{\rm t}$ and $\beta$. The points are colored by accretion rate (Equation 4). The cyan squares are from the $\Gamma$-binned median profiles. The black dashed line is a linear fit.
@@
In Figure 5 we plot the accretion rate dependence of the stellar $r_{\rm t}$ and $r_{\rm steep}$. For comparison, we also show the corresponding plots for the dark matter. The light blue points are individual clusters, while the dark blue points show the median in the same accretion rate bins as we use to calculate median profiles. The measurements from the median profiles are shown in red squares. The increased transparency on the right-most points indicates that they correspond to bins with <20<20 clusters. The fitting function for $\Gamma_{200{\rm m}}-r_{\rm steep}$ from More et al. (2015) is plotted on the top row and agrees well with our data. The fitting function for $\Gamma_{200{\rm m}}-r_{\rm t}$ from Diemer (2025) is plotted on the bottom row and agrees at large $\Gamma_{200{\rm m}}$, but rises more rapidly than our data towards the small end. For comparison, we plot the More et al. (2015) relation rescaled according to the $r_{\rm t}$-$r_{\rm steep}$ relation we find from Figure 4, and it seems to be a better fit.

### v2

@@ Abstract
pubyear: 2026pagerange: The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profilesB
@@ 1 Introduction
The precise definition of a galaxy cluster’s outer boundary is fundamental to studies of cluster formation, evolution, and cosmology (e.g. Kravtsov and Borgani, 2012; Morandi and Sun, 2016; Haggar et al., 2024). Traditional definitions—deduced from friends-of-friends (FoF) linking lengths or spherical overdensity radii—play a key role in determining cluster masses and membership (e.g. Eke et al., 2004; Lukić et al., 2009), yet these are choices driven primarily by what is practicable with data rather than what is most physically meaningful (e.g. White, 2001; Old et al., 2014, 2015). Establishing a boundary that is physically meaningful and practicably measurable via observations remains a major challenge.
@@
A promising physically motivated boundary is provided by the splashback radius, which marks the apocentre of recently accreted material orbiting within the cluster potential (e.g. Diemer and Kravtsov 2014; Adhikari et al. 2014; More et al. 2015; Diemer 2022). This radius delineates the transition between orbiting and infalling matter, and manifests as a steepening in the halo density profile (which, we note, has historically—and erroneously—been conflated with the splashback radius itself). The radius of this steepest slope is of particular interest because there is now considerable evidence that this steepening has been measured in observations of real galaxy clusters (e.g. More et al. 2016; Baxter et al. 2017; Nishizawa et al. 2018; Chang et al. 2018b; Contigiani et al. 2019; Xu et al. 2024). This is exciting because there is good reason to expect that splashback radius offers a direct probe of halo assembly history and mass accretion rates, thereby connecting cluster outskirts to cosmological growth (More et al., 2015; Deason et al., 2021; Dacunha et al., 2025).
Observational detections of splashback features have been reported using satellite galaxy counts (e.g. More et al., 2016; Baxter et al., 2017; Nishizawa et al., 2018; Adhikari et al., 2021) and weak gravitational lensing (e.g. Chang et al., 2018b; Contigiani et al., 2019; Xu et al., 2024; Giocoli et al., 2024). Initial observations suggested that the splashback radius is smaller than that predicted by $\Lambda{\rm CDM}$ simulations by about $20\%$ (More et al., 2016; Baxter et al., 2017; Chang et al., 2018b). It has since been argued that this is the result of biases arising from projection effects in optically-selected clusters (Busch and White, 2017; Sunayama et al., 2020), particularly when using apertures significantly smaller than the splashback radius (Murata et al., 2020). These biases have been mitigated by better optical identification (Murata et al., 2020), or avoided altogether by selecting clusters by their Sunyaev-Zel’dovich (SZ) signal (Shin et al., 2019; Zürcher and More, 2019; Adhikari et al., 2021) or X-ray luminosity (Rana et al., 2023; Joshi et al., 2026).
@@
Methods for detecting the steepening feature have their own limitations. In the case of satellite galaxy counts, accurate measurement of splashback features depends on the properties of satellite tracers in a given cluster (e.g. O’Neil et al., 2022), and accounting for the complexity of their kinematics (e.g. Aung et al., 2021) and structure of the local cosmic web (e.g. Lebeau et al., 2024; Sun et al., 2025). Satellite galaxies are also biased tracers of the cluster potential, with the most massive tending to accumulate towards the central regions due to dynamical friction (Ludlow et al., 2009; Adhikari et al., 2016). Weak lensing directly probes the total projected mass profile and is free of such biases, but its signal-to-noise ratio is typically low because individual systems contain only a small number of lensed background sources.
@@
A promising alternative probe is the diffuse intra-cluster light (ICL) that is formed from the remnants of disrupted galaxies. This material is a potentially powerful tracer of cluster outskirts (Montes and Trujillo, 2019; Contini, 2021; Contreras-Santos et al., 2024), and a potential detection of the steepening feature in the ICL of an individual cluster has already been reported by Gonzalez et al. (2021). Theoretical work by Deason et al. (2021) has shown a direct correspondence between the steepening feature in the total stellar and dark matter density, indicating a dynamical link between these components. However, there remain important questions about the detailed nature of the stellar splashback structure, its relation to the underlying dark matter, and its observational accessibility, which we seek to answer in this paper.
@@
A deeper understanding of the stellar splashback feature is crucial for a number of reasons. The distribution of stars in a cluster reflect physical processes that affect satellite galaxies and their dark matter halos, such as tidal stripping (Sifón et al., 2018; Kumar and More, 2026; Xie and Gao, 2015) and dynamical friction (Ludlow et al., 2009; Adhikari et al., 2016), and so their distribution encodes information about the interplay between baryons and dark matter in the cluster outskirts (e.g. Brown et al., 2024). When combined with splashback and shock structures in the stars and gas (e.g. Deason et al., 2021; Towler et al., 2024; Zhang et al., 2025), this information can be used to infer dynamical history. Similarly, projected stellar density profiles provide a complementary means of constraining cluster growth histories (e.g. Caminha et al., 2017; Montes and Trujillo, 2018; Kluge et al., 2025; Joshi et al., 2026). This is especially valuable when supplemented by weak lensing; combining different datasets, affected by projection in different ways, allows for a more robust inference of splashback features (e.g. Adhikari et al., 2021).
@@
In this paper, we build upon recent work by Diemer (2022, 2023, 2025) and investigate the splashback region surrounding galaxy clusters by distinguishing material that is orbiting within the halo from material that is infalling for the first time. To do this, we use the statistical sample of massive galaxy clusters from The Three Hundred collaboration’s hydrodynamical galaxy formation simulation suite (cf. Cui et al., 2018, 2022). This is a mass complete sample of clusters drawn from a 1 $h^{-1}$ Gpc box, which have a diversity of assembly histories and larger-scale environments. These simulations self-consistently model the physics of galaxy formation, including star formation, black hole growth, supernovae, and AGN feedback. By investigating the stellar splashback structure and its relation to dark matter in hydrodynamical simulations, our aim is to establish a physically grounded framework for interpreting stellar density profiles as tracers of cluster assembly. This approach promises to enhance the utility of stellar observables—such as the ICL—in constraining cluster boundaries and mass accretion histories, thereby broadening the toolkit available for cluster cosmology.
@@ 2 The data
For this work we use data from the gizmo-simba run of the Three Hundred Project (Cui et al., 2018), a suite of hydrodynamical re-simulations of the most massive 324 galaxy cluster halos in the dark matter-only MultiDark Planck 2 (MDPL2) simulation (cf. Klypin et al., 2016), a $1h^{-1}$ Gpc box on a side. These clusters have virial masses in the range $6.4\times 10^{14}h^{-1}M_{\odot}\lesssim M_{200{\rm c}}\lesssim 2.6\times 10^{15}h^{-1}M_{\odot}$, where $M_{200{\rm c}}$ is the mass corresponding to an overdensity criterion of 200 times the critical density at that epoch. Each re-simulated region extends $15\,h^{-1}\,{\rm Mpc}$ in comoving radius from the $z=0$ cluster centre, corresponding to several virial radii.
The re-simulations were run using the meshless finite mass (MFM) implementation of the gizmo hydrodynamics code (cf. Hopkins, 2015). This method overcomes some of the limitations of classic smoothed-particle hydrodynamics and moving-mesh approaches and captures fluid mixing and instabilities. The gravity solver used in gizmo is derived from gadget-3 (Springel, 2005). Galaxy formation physics is implemented using a modified version of the simba model (Davé et al., 2019) that has been calibrated for cluster scales (cf. Cui et al., 2022). It models radiative cooling, star formation and feedback, black hole formation and growth, and multiple modes of black hole feedback. Star particles are spawned from gas elements stochastically, with each MFM fluid element spawning a single star particle of the same mass. The gizmo-simba model was run on re-simulation regions from both a $3840^{3}$- and a $7680^{3}$-particle box with initial conditions from MDPL2, which we refer to as GIZMO_3k and GIZMO_7k, respectively. Unless stated otherwise, in this work we present results from the GIZMO_7k run, but we have verified convergence using results from the GIZMO_3k.
Particles inside the re-simulation regions were split into dark matter and gas particles with masses of $15.9\times 10^{7}\,h^{-1}\,{\rm M}_{\odot}$ and $2.95\times 10^{7}\,h^{-1}\,{\rm M}_{\odot}$, respectively, in the GIZMO_7k run, and eight times larger in the GIZMO_3k run. The Plummer-equivalent force softening used in the GIZMO_7k run is $\epsilon_{\rm P}=2.5\,h^{-1}\,{\rm kpc}$, which corresponds to a spline softening length as used by gadget-3 of $\epsilon_{\rm sp}=2.8\times\epsilon_{\rm P}=7\,h^{-1}\,{\rm kpc}$. When fitting radial profiles, we adopt a minimum radius equal to the halo convergence radius, which varies from halo to halo, but is always larger than $\epsilon_{\rm P}$ (Power et al., 2003; Ludlow et al., 2019).
@@
Data are saved in 129 snapshots from $z=16.98$ to $z=0$. For each cluster, we use group catalogues constructed with the AHF halo finder (cf. Gill et al., 2004; Knollmann and Knebe, 2009), which includes information about the stellar and gas content of the main halo and its substructures.
@@ 3.1 Orbiting-infalling decomposition
Figure 1: Cluster 001 from the GIZMO_3k run, decomposed into orbiting (yellow-purple) and infalling (black) particles. The top row shows the xxyy projection of the dark matter (left) and stars (right), while the bottom row shows their radial phase space. The cluster’s $R_{200{\rm c}}$ and $R_{200{\rm m}}$ are marked.
@@
Figure 2: The density profiles of the orbiting and infalling components, fit using Equations 13. The BCG component is confined to the inner regions, while the stellar “halo” takes on a similar shape to the orbiting profile of the dark matter. The logarithmic slope profiles are shown in the bottom panels.
The cluster center of mass and velocity center are calculated using particles of all types (dark matter, stars, and gas) that are contained within $R_{200{\rm c}}$ of the position of peak density determined by ahf. To identify the splashback of stars and dark matter in the simulated clusters, we classify all particles as “orbiting” or “infalling” depending on if the particle has passed through at least one pericenter about the center of mass of the cluster. This is done by recording a change in the sign of a particle’s physical radial velocity (peculiar plus Hubble flow) from negative to positive. To filter out radial velocity sign flips due to orbits within subhalos, we also require that particles have traversed an angle of at least $\pi/2$ around the halo center since entering the cluster region. Any particle that has not had such a sign change is regarded as “infalling”. Though this classification technically considers positive radial velocity particles beyond the turn-around radius as “infalling”, we limit our analysis of the density profiles to radii safely within this boundary. We also do not track gas particles, meaning any stars that form in orbit are temporarily classified as “infalling” until their first pericenter.
We identify cluster progenitors by selecting the 100 most-bound particles of each halo and identifying the halo in the previous snapshot that contains the highest fraction of them. This gives us a halo position $\boldsymbol{X}(z)$ and velocity $\boldsymbol{V}(z)$ at every snapshot relative to which we calculate the radial velocity of every particle. That is, for a particle with comoving position $\boldsymbol{x}$ and peculiar velocity $\boldsymbol{v}=a\dot{\boldsymbol{x}}$ (where $a$ is the scale factor) we calculate $v_{r}=[(\boldsymbol{v}-\boldsymbol{V})+H\boldsymbol{r}]\cdot\hat{\boldsymbol{r}}$, where $\boldsymbol{r}=a(\boldsymbol{x}-\boldsymbol{X})$.
The result of the orbiting-infalling decomposition for the most massive cluster in the GIZMO_3k run is shown in Figure 1. The top row shows the xxyy projection of the dark matter (left) and stars (right), while the bottom row is their radial phase space (with Hubble flow added). The phase space shows a clear separation between the infalling stream and the orbiting material, with the infalling material abruptly transitioning into orbiting material at the point of pericenter, $v_{\rm r}=0$. The orbiting particles at the largest radii still have some positive radial velocity, meaning they are still moving outwards and have not quite “splashed back”. Visually, the distribution of the dark matter is well traced by the stellar distribution, and both the orbiting dark matter and orbiting stars extend significantly past $R_{200{\rm c}}$ and out to about $R_{200{\rm m}}$. See e.g. Diemer (2017); Bakels et al. (2021) for discussion on the extent of the orbits of halo material.
@@ 3.3 Profile fitting functions
We fit the profiles using least-squares minimization in the logarithm of the density. The radial bins are weighted equally and the fitting is done between $4\,R_{200{\rm m}}$ and the convergence radius predicted by Power et al. (2003) and Ludlow et al. (2019) for the dark matter-only case. The results of fitting Equations 13 to the orbiting and infalling profiles of an individual cluster (cluster 001) is shown in Figure 2.
@@
where $a$ is the scale factor and $\Delta$ represents a change over one dynamical time, defined as the crossing time $2\,R_{200{\rm m}}/V_{200{\rm m}}$ $\propto(G\rho_{\rm m})^{-1/2}$. At $z=0$, one dynamical time in the past corresponds to $a\approx 0.769$. Here, as with everywhere else in this work, $M_{200{\rm m}}$ is calculated using the total density (stars, gas, and dark matter). Progenitor halos are identified as those that contain the largest fraction of the most-bound particles of the $z=0$ halos. We choose six equal-width bins between $\Gamma=0$ and $\Gamma=6$ and calculate the median logarithmic density profile in each bin, in units of $r/R_{200{\rm m}}$. The first four bins contain between 51 and 102 clusters, while the remaining two contain 17 and 10.
@@ 4 Results
Figure 3: Comparison between $r_{\rm t}$ and $\beta$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The square points are the result of fitting the $\Gamma$-binned median profiles.
@@ 4.1 Fit parameter relations
Figure 4: Comparison between the radius of steepest slope, $r_{\rm steep}$ and $r_{\rm t}$ for the dark matter and stars. The points are colored by accretion rate (Equation 4). The square points are the result of fitting the $\Gamma$-binned median profiles. The black dashed lines are linear fits given by $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.41$ for the dark matter and $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.37$ for the stars.
Figure 5: $r_{\rm steep}$ and $r_{\rm t}$ as a function of accretion rate (Equation 4). The light blue points are measurements for individual profiles, while the dark blue points are the median and standard deviation in equally-spaced bins in $\Gamma$. The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), while in the bottom row it is the same relation rescaled using the best-fit $r_{\rm t}$$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025).
@@
The truncation radius is a more direct measure of the size of the splashback boundary (technically the orbiting boundary, since the outermost particles may not have reached apocenter yet) than the radius of steepest slope, which is determined in part by the infalling material. In perfectly spherical collapse, the truncation of the orbiting profile is infinitely sharp and the two radii coincide. However, in realistic simulations, mass is accreted anisotropically and particle orbits can take on a variety of shapes, which has the effect of smoothing out the edge of the orbiting profile. In this case, the radius of steepest slope is not simply related to the edge of the orbiting material, but is rather the point at which the infalling profile becomes dominant enough to strongly influence the total slope. In Figure 4 we plot $r_{\rm steep}$ measured from the total profiles against $r_{\rm t}$. Again, the points are coloured by $\Gamma$ and show the gradient expected from the now well-known $r_{\rm steep}$$\Gamma$ relation. The cyan points are measurements of $r_{\rm steep}$ from the $\Gamma$-binned median total profile. The $r_{\rm t}$$r_{\rm steep}$ relation seems to be described well by a line with slope of $\sim 0.83$ and $r_{\rm steep}(r_{\rm t}=0)/R_{200{\rm m}}\approx 0.41$ for the dark matter and $0.37$ for the stars (black dashed lines). On average, the radius of steepest slope is $\sim 0.25\,R_{200{\rm m}}$ larger than the truncation radius.
@@
The gradient with accretion rate seen in the radii in Figures 3 and 4 reflect the now well-established relationship between the splashback radius and accretion rate. This relationship was first reported by Diemer and Kravtsov (2014) for the radius of steepest slope in dark matter-only cosmological simulations, while its theoretical basis was established by Adhikari et al. (2014) using a model based on spherical collapse. The relation was shown by More et al. (2015) to be well-described by an exponential of the form $r_{\rm steep}/R_{200{\rm m}}=a+be^{-\Gamma/c}$, and this same function was shown by Deason et al. (2021) to hold for the stellar $r_{\rm steep}$ of 34 C-EAGLE clusters. Recently, Diemer (2025) showed that the relationship between $r_{\rm t}$ and $\Gamma$ for stacked halos is also well-described by an exponential that has remarkably little dependence on halo mass or cosmology.
@@
Figure 7: Comparison between 3D and projected $r_{\rm t}$ and $\beta$. The points are colored by accretion rate (Equation 4). The square points are the result of fitting the $\Gamma$-binned median profiles. The black dashed line is a linear fit.
@@
In Figure 5 we plot the accretion rate dependence of the stellar $r_{\rm t}$ and $r_{\rm steep}$. For comparison, we also show the corresponding plots for the dark matter. The light blue points are individual clusters, while the dark blue points show the median in the same accretion rate bins as we use to calculate median profiles. The measurements from the median profiles are shown in red squares. The increased transparency on the right-most points indicates that they correspond to bins with <20<20 clusters. The fitting function for $\Gamma_{200{\rm m}}$$r_{\rm steep}$ from More et al. (2015) is plotted on the top row and agrees well with our data. The fitting function for $\Gamma_{200{\rm m}}$$r_{\rm t}$ from Diemer (2025) is plotted on the bottom row and agrees at large $\Gamma_{200{\rm m}}$, but rises more rapidly than our data towards the small end. For comparison, we plot the More et al. (2015) relation rescaled according to the $r_{\rm t}$$r_{\rm steep}$ relation we find from Figure 4, and it seems to be a better fit.